OpenStudy (anonymous):

solve: 6(x+4)=3(4x-2)

OpenStudy (aa1602):

6(x+4) = 3 ( 4x-2) 6x +24 = 12x - 6 6x - 12x = -6-24 -6x = -30 / x = 5